OpenStudy (natsent06):

Subtract 5y^2 + 2y - 8 from 4y^2 - 5y + 1.

OpenStudy (misty1212):

\[4y^4-5y+1-(5y^2+2y-8)\] is how to set up the problem because in english we say it backwards

OpenStudy (misty1212):

get rid of the parentheses by distributing the minus sign, then combine like terms

OpenStudy (natsent06):

there were no parenthesis though n the problem

OpenStudy (misty1212):

yes i know there were not , but is said "subtract" which means subtract the whole thing

OpenStudy (natsent06):

oh

OpenStudy (misty1212):

like if you said "subtract 24 from 59" it would be \(59-24\) you would subtract 4 from 9 and 2 from 5, subtract the whole thing

OpenStudy (phi):

can you finish? so far you have \[ 4y^2-5y+1-(5y^2+2y-8) \]

OpenStudy (phi):

fixed typos. so far you have \[ 4y^2-5y+1+ -1\cdot 5y^2 + -1\cdot 2y + -1 \cdot -8 \]

OpenStudy (natsent06):

i think i figured it out but i have more questions i dont get

OpenStudy (phi):

what did you get ?

OpenStudy (natsent06):

uh i changed questions allready on my quiz